The surface area A of an average human body is established to be given by A(w,h)=16.54w0.75h0.25 where w is the weight and h is the height.

  1. Find AW(65,57) and Ah(65,57), interpret the results.
  2. For a 67 pounds child whose height is 57 inches determine the change in area A if weight decreases by 5 pounds and height increases by3 inches.
1
Expert's answer
2021-07-27T15:24:23-0400

"A(w,h)=16.54w^{0.75}h^{0.25}"


"A_w(w,h)=\frac{\partial A}{\partial w}= 16.54\cdot 0.75w^{-0.25}h^{0.25}=12.405(h/w)^{1/4}"

"A_w(65,57)=12.405(57\/65)^{1\/4}=12.0043"

This equality means that the surface area A of an average human body with weight and heght close to (65, 57) changes by approximately 12.0043 per each pound.


"A_h(w,h)=\frac{\partial A}{\partial h}= 16.54\cdot 0.25w^{0.75}h^{-0.75}=4.135(h/w)^{-3/4}"

"A_h(65,57)=4.135(57\/65)^{-3\/4}=4.653"

This equality means that the surface area A of an average human body with weight and heght close to (65, 57) changes by approximately 4.653 per each inch.


If "w=65", "h=57", "\Delta w=5" and "\Delta h=3", then the change in area A will be approximately equal to

"\Delta A\approx A_w\Delta w + A_h\Delta h = 12.0043\cdot 5+4.653\cdot3=72.0258"